(a) How many code words over a, b, c, d of length 20 contain exactly 10 a’s?

(b) How many contain exactly 10 a’s and 5b’s.


1
Expert's answer
2021-05-18T13:54:10-0400

(a) Remaining 20-10=10 places to be filled with b,c,d. Each of these places can be filled in 3 ways. Hence total number of ways of filling is "3^{10}." Out of 20 places 10 places can be chosen in "(\begin{matrix}
20\\
10
\end{matrix})" ways. Hence answer is "3^{10}""(\begin{matrix}
20\\
10
\end{matrix})" .

(b) Here as before we need to fill 5 places with c and d. This can be done in "2^5" ways. The 5 places can be chosen in "(\begin{matrix}
20\\
5
\end{matrix})" ways. Now a and b with first one 10 times repeated and b 5 times repeated can be done in "\frac{15!}{10!5!}" ways. Hence no of words is "2^5 (\begin{matrix}
20\\
5
\end{matrix})" "\frac{15!}{10!5!}" ways.